OpenStudy (calculusxy):

MEDAL for help!!!! Question is posted below:

OpenStudy (calculusxy):

Write a set of walking instructions based on the data in Tables A and B.

OpenStudy (calculusxy):

@Michele_Laino

OpenStudy (calculusxy):

What do they mean by "walking instructions''?

OpenStudy (michele_laino):

maybe the function distance as function of time for both tables A and B

OpenStudy (michele_laino):

I'm not sure

OpenStudy (michele_laino):

do you know the least square method, for statistical processing of data?

OpenStudy (calculusxy):

No

OpenStudy (calculusxy):

But I can learn !

OpenStudy (michele_laino):

what subject are you studying at your school?

OpenStudy (calculusxy):

Well i am a rising eight grader...

OpenStudy (calculusxy):

And it's just math, latin, english, history, and physics.

OpenStudy (michele_laino):

please wait, I'm working on your answer...

OpenStudy (calculusxy):

Okay

OpenStudy (michele_laino):

I understand. We have to write two functions which model your data contained in both tables

OpenStudy (michele_laino):

for example, let's consider the first table

OpenStudy (calculusxy):

so can we use: y = mx +b

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (calculusxy):

Okay so to find m: \[\frac{ (y_2 - y_1) }{ (x_2 -x_1) }\]

OpenStudy (michele_laino):

when x=0, we have y=0.8, so we can write: 0.8=m*0+b=b so b=0.8

OpenStudy (calculusxy):

So i am going to use the coordinates: (1,1) and (2,1.2) \[\frac{ (1.2 - 1) }{ (2 -1 ) }\]

OpenStudy (michele_laino):

similarly, when x=1, we have y=1 so we can write: 1=m*1+b=m+b=m+0.8 then: m=1-0.8=0.2 I prefer the substitution method

OpenStudy (calculusxy):

\[\frac{ 0.2 }{ 1 }\]

OpenStudy (michele_laino):

ok! so we can write: y=0.2*x+0.8

OpenStudy (calculusxy):

yes

OpenStudy (michele_laino):

now we have to do the same, with the second table

OpenStudy (calculusxy):

for b:\[\frac{ 3.2 - 3.6 }{ 3 -2 }\]

OpenStudy (calculusxy):

That would show a negative slope right?

OpenStudy (michele_laino):

yes! it is -0.4

OpenStudy (calculusxy):

\[\frac{ -0.4 }{ 1 }\]

OpenStudy (michele_laino):

ok!

OpenStudy (calculusxy):

so y = -0.4x + 4

OpenStudy (michele_laino):

so the requested function is: y=-0.4*x+4

OpenStudy (calculusxy):

So those are the two answers right?

OpenStudy (michele_laino):

yes! right!

OpenStudy (calculusxy):

Can you please help me with my physics homework as well?

OpenStudy (michele_laino):

ok!

OpenStudy (calculusxy):

Thank you i will make a new post

OpenStudy (michele_laino):

ok!